OpenStudy (anonymous):

Find the solution to this system of equations.. 8x - 4y = -32 -3x + 4y = 42

OpenStudy (anonymous):

Do you have any options for x and y?

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

No I dont'.. Sorry was eating.

OpenStudy (anonymous):

That's cool.

OpenStudy (anonymous):

y=-20 and x=34. That's what I got but I not sure it's right. Been a while since I've worked these.

OpenStudy (anonymous):

Do me a favor and close this question.

OpenStudy (anonymous):

@texaschic101 Can you verify?

OpenStudy (texaschic101):

8x - 4y = -32 -3x + 4y = 42 --------------add 5x = 10 x = 2 -3x + 4y = 42 -3(2) + 4y = 42 -6 + 4y = 42 4y = 42 + 6 4y = 48 y = 12 check.. 8x - 4y = -32 8(2) - 4(12) = -32 16 - 48 = -32 -32 = -32 (correct) x = 2 and y = 12

OpenStudy (anonymous):

THANK YOU SO MUCH. YOU'RE A GOD IN MY EYES OF MATHEMATICS XD.

OpenStudy (texaschic101):

:) lol...glad to help